### v4

# DAMR: Efficient and Adaptive Context-Aware Knowledge Graph Question Answering with LLM-Guided MCTS

Abstract
Knowledge Graph Question Answering (KGQA) aims to interpret natural language queries and perform structured reasoning over knowledge graphs by leveraging their relational and semantic structures to retrieve accurate answers. Existing methods primarily follow either the retrieve-then-reason paradigm, which relies on Graph Neural Networks (GNNs) or heuristic rules to extract static candidate paths, or dynamic path generation strategies that employ large language models (LLMs) with prompting to jointly perform retrieval and reasoning. However, the former lacks adaptability due to static path extraction and the absence of contextual refinement, while the latter suffers from high computational costs and limited evaluation accuracy because of their dependence on fixed scoring functions and repeated LLM calls. To address these issues, this paper proposes Dynamically Adaptive MCTS-based Reasoning (DAMR), a novel framework that integrates LLM-guided Monte Carlo Tree Search (MCTS) with adaptive path evaluation to enable efficient and context-aware KGQA. DAMR leverages MCTS as a backbone, where an LLM-based planner selects the top-$k$ semantically relevant relations at each expansion step to effectively reduce the search space. To enhance evaluation accuracy, we introduce a lightweight Transformer-based scorer that performs context-aware plausibility estimation by jointly encoding the question and relation sequence through cross-attention, thereby capturing fine-grained semantic shifts during multi-hop reasoning. Furthermore, to mitigate the scarcity of high-quality supervision, DAMR incorporates a dynamic pseudo-path refinement mechanism that periodically generates training signals from partial paths explored during search, enabling the scorer to continually adapt to the evolving distribution of reasoning trajectories. Extensive experiments on multiple KGQA benchmarks show that DAMR significantly outperforms state-of-the-art methods.
1 Introduction
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities across diverse tasks, including mathematical problem solving (pei2025mathfusion, ; didolkar2024metacognitive, ), commonsense inference (toroghi2024verifiable, ), and open-domain question answering (zhao2023divknowqa, ). Despite their generalization ability, LLMs often struggle in domain-specific scenarios due to the lack of grounded external knowledge, resulting in factual hallucinations and high inference costs (huang2025survey, ; wang2024factuality, ). To address these limitations, recent efforts have explored integrating domain knowledge into LLM reasoning. A promising direction to overcome these limitations is Knowledge Graph Question Answering (KGQA) (dammu2025dynamic, ; saxena2020improving, ; choi2023nutrea, ), which integrates relational structures into the reasoning process to provide factual grounding and structural interpretability. By combining the expressiveness of natural language with the precision of knowledge graphs, KGQA offers a scalable solution to improve factual consistency, reasoning transparency, and answer reliability (liu2025dual, ; yao2025learning, ).
Existing KGQA approaches can be broadly categorized into two paradigms according to how reasoning paths are constructed: retrieve-then-reason methods and dynamic path generation strategies. In the former, candidate paths are extracted prior to answer prediction, typically using Graph Neural Networks (GNNs) (ma2025debate, ; yao2025learning, ) or rule-based heuristics (fang2024karpa, ). However, such methods exhibit limited adaptability, as GNNs fail to incorporate question-specific semantics during inference, while heuristic rules remain inherently inflexible and cannot support dynamic refinement of reasoning (liu2025dual, ; yao2025learning, ). In contrast, dynamic path generation strategies unify retrieval and reasoning by constructing paths dynamically during question processing. These methods either leverage LLMs to iteratively generate paths through in-context learning or Chain-of-Thought (CoT) (sui2024fidelis, ; li2024framework, ), or employ guided search techniques such as MCTS, where paths are incrementally expanded with a path scorer (ma2025deliberation, ; shen2025reasoning, ). While offering greater flexibility, such approaches suffer from substantial computational overhead due to repeated LLM calls and limited evaluation accuracy, as static scorers cannot capture the evolving semantics of reasoning paths (chang2024survey, ; lee2024zero, ).
This paper investigates the design of an adaptive KGQA framework to address challenges of computational inefficiency and limited path evaluation accuracy in dynamic reasoning. However, developing such a framework raises three key challenges: (1) How to modularize reasoning to reduce LLM overuse in search? Inefficiency in dynamic KGQA stems from repeatedly invoking LLMs for relation retrieval and reasoning in multi-hop path construction (shen2025reasoning, ; long2025enhancing, ). Although methods such as CoT and MCTS enable flexible exploration, they bind LLMs to every decision step, causing high inference cost and poor scalability. The challenge is to design a modular framework that leverages LLMs efficiently, guiding search without direct involvement at each step. (2) How to accurately evaluate evolving reasoning paths? As multi-hop paths are incrementally constructed, their semantics evolve with each added relation and context. Yet existing methods rely on static scoring or shallow similarity metrics that fail to capture these semantic shifts (xu2024llm, ; sui2024fidelis, ). This highlights the challenge of designing a path evaluation model that adaptively captures fine-grained changes conditioned on both the question and evolving relation sequence. (3) How to train a reliable evaluation model with limited supervision? Accurate path ranking requires a well-calibrated scorer, yet dynamic reasoning generates many incomplete paths with only a few valid ones. This results in imbalanced, noisy supervision, especially for multi-hop questions where successful trajectories are scarce. While reinforcement learning has been explored (ma2024coevolving, ; zhai2024fine, ), it suffers from sparse rewards and unstable optimization. The key challenge is to construct reliable learning signals from limited supervision to support adaptive scorer training.
To address these challenges, we propose Dynamically Adaptive MCTS-based Reasoning (DAMR), an efficient framework that integrates LLM-guided MCTS with context-aware semantic modeling for accurate and efficient KGQA. DAMR employs an MCTS backbone, where an LLM-based planner dynamically guides path expansion by selecting semantically relevant relations at each step, thereby reducing the search space and improving answer identification. For path evaluation, we introduce a lightweight Transformer-based scorer that jointly encodes the question and relation sequence via cross-attention, enabling context-sensitive plausibility estimation and capturing evolving semantics during multi-hop reasoning. To mitigate supervision scarcity, DAMR further incorporates a dynamic pseudo-path mechanism that continuously adapts the scorer during search: partial paths from MCTS rollouts are ranked by predicted plausibility and converted into pseudo-path supervision pairs, amplifying signals from promising trajectories while suppressing noise from suboptimal ones. Our contributions are summarized as follows:
  • We study adaptive path reasoning in KGQA, where the key challenges lie in capturing the evolving semantics of multi-hop reasoning paths and ensuring computational efficiency during search, motivating the need for dynamic and context-aware reasoning strategies.
  • We propose DAMR, a novel framework that integrates MCTS with a dynamically adapted path evaluation model, enhancing evaluation accuracy while maintaining computational efficiency.
  • We conduct extensive experiments across multiple KGQA benchmarks, demonstrating that DAMR consistently outperforms state-of-the-art methods.
2 Related work
Knowledge Graph Question Answering (KGQA).
KGQA aims to enhance reasoning capabilities by incorporating external knowledge graphs to answer natural language questions (choi2023nutrea, ; xu2025memory, ). Existing KGQA approaches can be broadly classified into two categories: retrieve-then-reason and dynamic path generation. The first category extracts candidate reasoning paths using Graph Neural Networks (GNNs)(ma2025debate, ; yao2025learning, ) or rule-based heuristics(fang2024karpa, ), followed by LLM-based answer generation. While GNNs learn embeddings to identify relevant paths and rule-based methods apply predefined patterns (zhao2023divknowqa, ; liu2025dual, ), these approaches lack the flexibility to adapt dynamically to question-specific context during inference. In contrast, dynamic path generation methods, such as CoT prompting (sui2024fidelis, ; li2024framework, ) and MCTS (ma2025deliberation, ; shen2025reasoning, ), unify retrieval and reasoning for more flexible exploration. However, they suffer from high computational overhead due to repeated LLM calls, and static scorers often fail to adapt to evolving path semantics (long2025enhancing, ; luo2025kbqa, ; ma2024think, ). To address these challenges, we propose an adaptive framework that integrates symbolic search with a fine-tuned evaluation model, aiming to improve both computational efficiency and reasoning accuracy in KGQA.
Adaptive and Self-Improving Reasoning Models. A promising approach to developing adaptive reasoning models is to frame the process within a reinforcement learning (RL) paradigm, where an agent learns a policy to navigate a state space. Early methods such as DeepPath (xiong2017deeppath, ) and MINERVA (das2018go, ) used RL to discover reasoning paths by rewarding the agent only when a correct answer is reached. However, this leads to the sparse rewards problem, as positive feedback arrives only after long action sequences, resulting in weak learning signals and poor exploration efficiency (zhai2024fine, ; chang2023learning, ). To address this challenge, an alternative is self-training via pseudo-labeling, where the model learns from its own high-confidence predictions (lee2013pseudo, ; xie2020self, ). While commonly used in semi-supervised learning, pseudo-labeling proves especially effective in reasoning tasks with limited supervision (wang2022semi, ; huang2025mllm, ). Instead of relying on sparse terminal rewards, we leverage intermediate search paths as dynamic pseudo-paths, offering dense and adaptive supervision. This facilitates continual refinement of the path evaluator to better capture the evolving semantics of reasoning.
3 Methodology
Figure 1: Overview of DAMR. The reasoning process begins with an MCTS guided by an LLM-based planner, which selects top-$k$ semantically relevant relations at each expansion step. A context-aware path evaluator scores each candidate path during simulation. To enable continual adaptation, high-confidence pseudo-paths generated during search are used to dynamically fine-tune the evaluator.
3.1 Problem Formulation
We define Knowledge Graph Question Answering (KGQA) as the task of answering a natural language question by reasoning over a knowledge graph (KG). The KG is typically represented as a set of triples $\mathcal{K}=\{(e_{s},r,e_{o})\}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and binary relations. The goal of KGQA is to find a set of answers $\mathcal{A}_{q}\subseteq\{(e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3})\cdots\}$ for question $q$, such that a reasoning path through the KG leads from a topic entity to the correct answer. Formally, this is often framed as mapping $q$ to an executable query program $p_{q}$, where $\text{LLM}(p_{q}|{\mathcal{K}})=\mathcal{A}_{q}$.
3.2 Overview of Framework
In this paper, we propose a dynamically adaptive reasoning framework DAMR for KGQA, as shown in Fig. 1. DAMR comprises three components: (1) LLM Guided Expansion. DAMR employs MCTS to incrementally expand reasoning paths, guided by an LLM-based planner that proposes relevant relations. This reduces computational overhead and enhances efficiency in KG exploration; (2) Context-Aware Path Evaluation. To capture the evolving semantics of reasoning paths, DAMR employs a lightweight Transformer-based scorer with cross-attention to jointly encode the question and path embeddings. This enables context-sensitive evaluation and improves the accuracy and relevance of multi-hop reasoning; (3) Path-based Dynamic Refinement. DAMR uses intermediate paths from MCTS as dynamic pseudo-paths to iteratively fine-tune the path evaluator, enhancing its ability to capture question-specific semantics and improving reasoning accuracy.
3.3 LLM Guided Expansion
A key challenge in KGQA is efficiently exploring the vast search space of multi-hop reasoning paths, especially under weak or no supervision. Existing methods often struggle to balance search efficiency and semantic relevance, resulting in either redundant exploration or missed correct paths (long2025enhancing, ; li2024simple, ). To address this, the LLM-Guided Expansion module employs MCTS (kocsis2006bandit, ) as the backbone for symbolic path expansion. At each step, an LLM proposes semantically relevant relations, narrowing the search space and improving path quality, while MCTS ensures a balanced trade-off between exploration and exploitation.
Specifically, each node in the MCTS represents a reasoning state anchored at a specific entity in the KG. Given the current state, possible actions correspond to selecting an outgoing relation to extend the reasoning path. During the Selection phase, the search traverses from the root to a leaf node by recursively choosing children with the highest Upper Confidence Bound for Trees (UCT) score (kocsis2006bandit, ), which is defined as:
$UCT=\frac{w_{i}}{n_{i}}+C\sqrt{\frac{\ln N}{n_{i}}},$ (1)
where $w_{i}$ denotes the accumulated reward of node $i$, $n_{i}$ its visit count, $N$ the visit count of its parent, and $C$ a constant that balances exploration and exploitation. This criterion guides the search to trade off between exploring new relations and exploiting promising paths. In the Expansion phase, the selected leaf node is expanded by retrieving the outgoing relations of its entity $e_{i}$, denoted as $\mathcal{R}_{e_{i}}=\{r_{1},r_{2},\dots,r_{n}\}$. To avoid exhaustive branching, we prompt an LLM with the question $q$ and candidate relations $\mathcal{R}_{e_{i}}$, selecting the top-$k$ relations most semantically aligned with $q$:
$\mathcal{R}_{top-k}=\text{LLM}(q,\mathcal{R}_{e_{i}}).$ (2)
These top-$k$ relations are then used to generate new child nodes, ensuring that expansion favors semantically meaningful directions and remains computationally efficient.
3.4 Context-Aware Path Evaluation
While LLM-guided expansion effectively narrows the search space by selecting semantically relevant relations, it does not ensure that all expanded paths remain valid in the broader reasoning context. As the search progresses, path semantics evolve dynamically, and early promising trajectories may later become misleading or irrelevant. Prior works attempt to filter or pre-plan paths to improve relevance (fang2024karpa, ; long2025eperm, ), but still leave gaps in capturing evolving path semantics during search. To address this, we integrate a lightweight Transformer-based path scorer into MCTS’s simulation phase, employing cross-attention to jointly encode the question and current path, enabling adaptive evaluation that reflects evolving semantics.
Context-Aware Path Evaluator. In the Simulation phase, we assess the quality of candidate paths generated during MCTS rollouts. Given a question $q$ and a candidate relation path $p_{r}=(r_{1},r_{2},\ldots,r_{l})$, where $p_{r}$ is formed by sequentially selecting relations during expansion, both $q$ and $p_{r}$ are first encoded using a pre-trained LLM. Let $\mathbf{z}_{q}\in\mathbb{R}^{d}$ denote the embedding of $q$ and $\mathbf{z}_{r_{i}}\in\mathbb{R}^{d}$ the embedding of relation $r_{i}$. To capture the sequential structure of $p_{r}$, we introduce a learnable positional encoding $\mathbf{e}_{i}^{\text{pos}}$ for each relation. The final input sequence is obtained by combining $\mathbf{z}_{r_{i}}$ with $\mathbf{e}_{i}^{\text{pos}}$ and feeding the sequence into a Transformer encoder:
$\mathbf{E}_{p_{r}}=\text{Transformer}([\mathbf{z}_{r_{1}}+\mathbf{e}_{1}^{\text{pos}},\ldots,\mathbf{z}_{r_{l}}+\mathbf{e}_{l}^{\text{pos}}]),$ (3)
where $\mathbf{e}_{i}^{\text{pos}}=\mathbf{E}^{\text{pos}}[i]$ is the positional embedding of the $i$-th hop, drawn from a trainable matrix $\mathbf{E}^{\text{pos}}\in\mathbb{R}^{L\times d}$, where $L$ is the maximum path length and $d$ the embedding dimension. To further incorporate question-specific information, we apply a cross-attention mechanism, allowing the encoded path representation $\mathbf{e}_{p_{r}}$ to attend to the question embedding $\mathbf{z}_{q}$:
$\mathbf{H}=\mathbf{E}_{p_{r}}+\mathrm{CrossAttn}(\mathbf{E}_{p_{r}},\mathbf{z}_{q}),\;with\;\text{CrossAttn}(\mathbf{E}_{p_{r}},\mathbf{z}_{q})=\text{softmax}({\mathbf{E}_{p_{r}}\cdot\mathbf{z}_{q}^{T}}/{\sqrt{d_{k}}})\cdot\mathbf{z}_{q}.$ (4)
We then apply attention pooling over the relation representations $\mathbf{H}$ to obtain the hidden state of the relation path:
$\mathbf{s}_{p_{r}}=\sum_{i=1}^{l}\alpha_{i}\textbf{h}_{i},\quad\alpha=\text{Softmax}\big(\text{MLP}(\mathbf{H})\big),$ (5)
where $\mathbf{h}_{i}$ is the hidden state of the $i$-th relation and $\alpha_{i}$ its learned attention weight. This pooling allows the model to emphasize informative steps along the reasoning path. The pooled path representation $\mathbf{s}_{p_{r}}$ is then concatenated with the question embedding $\mathbf{z}_{q}$, and the combined vector is passed through a multi-layer perceptron to compute the plausibility score of the question–path pair:
$S(q,p_{r})=\mathrm{MLP}\big([\mathbf{s}_{p_{r}}:\mathbf{z}_{q}]\big).$ (6)
This context-aware evaluator assigns plausibility scores to partial reasoning paths by jointly encoding the question and relation sequence, providing accurate, context-sensitive guidance for MCTS.
Pre-training of Evaluator. To train the context-aware evaluator, we construct supervision by sampling positive and negative paths from local subgraphs. A path is positive if it links the head entity to a correct answer within a hop limit. Negatives come from two sources: hard negatives that terminate near but miss the answer, and random negatives from walks that avoid answer entities. Each training instance is a triplet $(q,p^{+},p^{-})$, with sequences zero-padded and masked for efficient batch training.
The model computes a plausibility score $S(q,p)$ for each question-path pair and is optimized using the Pair-wise Ranking loss (rendle2012bpr, ) to encourage higher scores for positive paths:
$\mathcal{L}_{\text{PR}}=-\frac{1}{M}\sum_{i=1}^{M}\log\sigma\left(S(q,p^{+}_{i})-S(q,p^{-}_{i})\right),$ (7)
where $\sigma(\cdot)$ is the sigmoid function. This training strategy equips the evaluator with the ability to distinguish plausible reasoning paths, thereby improving the guidance signal during MCTS inference.
3.5 Path-based Dynamic Refinement
While LLM-guided expansion and semantic scoring improve path exploration, the static evaluator may fail to generalize to the evolving search space. To address this, we introduce a dynamic refinement mechanism that leverages high-confidence paths from MCTS rollouts as pseudo-paths. These pseudo-paths serve as supervision signals, enabling continual adaptation of the evaluator to new reasoning contexts without requiring additional labeled data.
Specifically, during Backpropagation phase, the plausibility score estimated by the context-aware path evaluator is propagated along the visited nodes in the MCTS tree after each simulation. For every entity $e_{i}$ on the simulated path, we update its visit count and aggregated value as follows:
$n_{e_{i}}=n_{e_{i}}+1,\quad w_{e_{i}}=\frac{\sum_{j}n_{e_{j}}\cdot w_{e_{j}}}{\sum_{j}n_{e_{j}}},$ (8)
where $n_{e_{i}}$ is the visit count and $w_{e_{i}}$ is the aggregated value of entity $e_{i}$. The value is computed as a weighted average over its child nodes $\{e_{j}\}$, and reflects the plausibility scores $w_{e_{j}}$ assigned during simulation. These updates refine the UCT estimates used in future selection steps, progressively biasing the search toward high-quality reasoning paths.
To construct supervision signals for fine-tuning, we dynamically sample pseudo-path pairs ($\hat{p}^{\prime}_{i}$, $\hat{p}^{\prime}_{j}$) from the set of explored paths during MCTS. Instead of relying on the evaluator’s predictions, we assign pseudo-labels based on empirically grounded values derived from the search process. Specifically, for entity $e_{i}$ along a reasoning path $p_{r}$, we define its search value as: $w_{e_{i}}=\frac{w_{p_{r}}}{n_{e_{i}}}$, where $w_{p_{r}}$ is the cumulative reward from all rollouts passing through $p_{r}$, and $n_{e_{i}}$ is the visite count of entity $e_{i}$. Given a pair of paths, we assign pseudo-labels based on their relative values:
$\left(\hat{p}^{+},\,\hat{p}^{-}\right)=\begin{cases}(p^{\prime}_{i},\,p^{\prime}_{j}),&\text{if }w_{e_{i}}>w_{e_{j}},\\ (p^{\prime}_{j},\,p^{\prime}_{i}),&\text{otherwise}.\end{cases}$ (9)
The path evaluator is then fine-tuned using the Pair-wise Ranking loss in Eq. 7, encouraging higher scores for more promising paths.
3.6 Reasoning Process
The overall reasoning process is outlined in Appendix A. The framework first initializes a path evaluator to discriminate between plausible and implausible KG paths, providing a foundation for downstream search. During dynamic MCTS, the algorithm iteratively performs selection, expansion, simulation, and backpropagation. In expansion, an LLM-based planner adaptively selects the top-$k$ relations most relevant to the question, steering the search toward semantically meaningful paths. The evaluator guides simulation by prioritizing trajectories likely to yield correct answers, while pseudo-path pairs sampled during search are periodically used for refinement. Finally, entities reached by high-scoring paths are aggregated to form the answer set.
4 Experiments
4.1 Experimental Settings
Datasets. To evaluate the effectiveness of DAMR, we conduct experiments on two widely used KGQA benchmarks: WebQSP talmor2018web and CWQ yih2016value . Following prior work sun2023think ; liu2025dual , we uniformly sample 1,000 questions from the test sets of both datasets to evaluate the performance. More details about datasets are provided in Appendix  B.
Baselines. We compare DAMR with a comprehensive set of baselines. These baselines include: the semantic parsing methods, e.g., KV-Mem (miller2016key, ), EmbedKGQA (saxena2020improving, ), QGG (lan2020query, ), NSM (he2021improving, ), TransferNet (shi2021transfernet, ), and KGT5 (saxena2022sequence, ); the retrieval-based methods, e.g., GraftNet (sun2018open, ), PullNet (sun2019pullnet, ), SR+NSM (zhang2022subgraph, ), and SR+NSM+E2E (zhang2022subgraph, ); the general LLMs, including Flan-T5-xl (chung2024scaling, ), Alpaca-7B (taori2023stanford, ), Llama3-8B (dubey2024llama, ), Qwen2.5-7B (team2024qwen2, ), ChatGPT (schulman2022chatgpt, ), and ChatGPT+CoT (wei2022chain, ); and recent LLMs with KG methods, including UniKGQA (jiang2022unikgqa, ),DECAF (yu2022decaf, ), KD-CoT (wang2023knowledge, ), Nutrea (choi2023nutrea, ), ToG (sun2023think, ), RoG (luo2023reasoning, ), KAPING (baek2023knowledge, ), ReasoningLM (jiang2023reasoninglm, ), FiDeLis (sui2024fidelis, ), GNN-RAG (mavromatis2024gnn, ), DoG (ma2025debate, ), DualR (liu2025dual, ) , DP (ma2025deliberation, ), and RwT (shen2025reasoning, ). The details are provided in Appendix C.
Type Methods WebQSP CWQ
Hits@1 F1 Hits@1 F1
Semantic Parsing KV-Mem (miller2016key, ) 46.7 34.5 18.4 15.7
EmbedKGQA (saxena2020improving, ) 66.6 - 45.9 -
QGG (lan2020query, ) 73.0 73.8 36.9 37.4
NSM (he2021improving, ) 68.7 62.8 47.6 42.4
TransferNet (shi2021transfernet, ) 71.4 - 48.6 -
KGT5 (saxena2022sequence, ) 56.1 - 36.5 -
Retrieval GraftNet (sun2018open, ) 66.4 60.4 36.8 32.7
PullNet (sun2019pullnet, ) 68.1 - 45.9 -
SR+NSM (zhang2022subgraph, ) 68.9 64.1 50.2 47.1
SR+NSM+E2E (zhang2022subgraph, ) 69.5 64.1 49.3 46.3
LLMs Flan-T5-xl chung2024scaling 31.0 - 14.7 -
Alpaca-7B taori2023stanford 51.8 - 27.4 -
Llama3-8B (dubey2024llama, ) 30.3 25.7 30.5 27.8
Qwen2.5-7B (team2024qwen2, ) 28.4 23.7 25.9 24.1
ChatGPT (schulman2022chatgpt, ) 66.8 - 39.9 -
ChatGPT+CoT (wei2022chain, ) 75.6 - 48.9 -
LLMs+KGs UniKGQA (jiang2022unikgqa, ) 77.2 72.2 51.2 49.0
DECAF yu2022decaf 82.1 78.8 70.4 -
KD-CoT wang2023knowledge 68.6 52.5 55.7 -
Nutrea (choi2023nutrea, ) 77.4 72.7 53.6 49.5
ToG (sun2023think, ) 81.9 76.0 68.5 60.2
RoG (luo2023reasoning, ) 80.8 70.8 57.8 56.2
KAPING (baek2023knowledge, ) 72.4 65.1 53.4 50.3
ReasoningLM (jiang2023reasoninglm, ) 78.5 71.0 69.0 64.0
FiDeLis (sui2024fidelis, ) 84.3 78.3 71.5 64.3
GNN-RAG (mavromatis2024gnn, ) 82.8 73.5 62.8 60.4
DoG (ma2025debate, ) 65.4 55.6 41.0 46.4
DualR (liu2025dual, ) 81.5 71.6 65.3 62.1
DP (ma2025deliberation, ) 87.5 81.4 75.8 69.4
RwT (shen2025reasoning, ) 87.0 79.7 72.4 66.7
DAMR 94.0 81.7 78.0 75.1
Table 1: Performance comparison (%) on WebQSP and CWQ datasets. Bold results indicate the best performance.
Implementation Details. We implement the DAMR framework using PyTorch, and all experiments are conducted on NVIDIA A100 GPUs. The LLM-based planner is implemented with GPT-4.1 (liu2023evaluating, ), while question and relation embeddings are generated from Qwen3-Embedding-8B (yang2025qwen3, ) with an embedding dimension of 1024. For the path evaluation module, we use a 128-dimensional embedding and employ the Adam optimizer (kingma2014adam, ) with a learning rate of $1\times 10^{-4}$ during pretraining and $1\times 10^{-5}$ during fine-tuning. The model consists of two Transformer layers and is trained for 15 epochs in the pretraining stage and 10 epochs in the fine-tuning stage. Following (luo2023reasoning, ; yao2025learning, ; ma2025deliberation, ), we evaluate DAMR using Hits@1 and F1 score, assessing answer correctness and overall accuracy for questions with potentially multiple correct answers.
4.2 Performance Comparison
We report the experimental results of DAMR in Table 1, benchmarking its performance against state-of-the-art baselines across KGQA datasets. From the results, we find that: (1) Semantic parsing and retrieval-based methods serve as early foundations for KGQA by extracting subgraphs and capturing structural semantics. However, embedding-based models struggle with complex relational patterns, while retrieval-based methods rely on rigid pipelines that limit generalization. In contrast, LLM with KG approaches combine the language understanding of LLMs with structured reasoning over KGs, enabling more flexible path exploration and improved adaptability to diverse, multi-hop queries. (2) General-purpose LLMs, such as ChatGPT and Llama3-8B, show basic reasoning ability but often perform worse than methods that combine LLMs with KGs in KGQA tasks. This is mainly because they are not grounded in domain-specific knowledge, making them more likely to produce incorrect or made-up answers. (3) DAMR consistently outperforms all baselines across both datasets, showcasing its strong reasoning capability. This superior performance is driven by its integration of an LLM-based planner, which selectively retrieves relevant relations to reduce noise and guide the search toward high-quality reasoning paths, and a path evaluation model that is dynamically fine-tuned during search to capture semantic differences among candidate paths and accurately rank those most likely to yield correct answers.
4.3 Efficiency Analysis
As shown in Table 3, DAMR achieves substantial improvements in computational efficiency. It reduces the average number of LLM calls to 7.1 on WebQSP and 16.8 on CWQ, with corresponding token usage of 3,931 and 9,266. These correspond to reductions of over 50% in LLM calls and 75% in token consumption relative to the strongest baseline. This efficiency is achieved by invoking the LLM only during the expansion phase of MCTS to select the top-$k$ semantically relevant relations, which effectively narrows the search space and avoids redundant reasoning steps that lead to unnecessary computational overhead. During simulation, the context-aware path evaluator efficiently assesses candidate paths based on question-path alignment without requiring any further LLM interaction or model inference. These design choices reduce both the frequency and verbosity of LLM usage while maintaining strong reasoning performance, making DAMR more efficient, scalable, and practically deployable than previous work.
4.4 Ablation Study
We conduct ablation studies to assess the contributions of key components in DAMR: (1) DAMR w/o PE: removing the path evaluation module; (2) DAMR w/o FT: disabling fine-tuning of the evaluator; and (3) DAMR w/ GPT-4.1: replacing the context-aware evaluator with a general LLM.
Experimental results are summarized in Table 3. From the results, we find that: (1) Removing the path evaluation module (DAMR w/o PE) causes a noticeable performance drop on both datasets, underscoring its critical role in guiding the search process. Without this component, the model cannot effectively assess or rank candidate paths, leading to suboptimal reasoning and reduced accuracy. (2) Compared to DAMR w/o FT, the proposed DAMR consistently achieves superior results on both datasets, highlighting the importance of the finetuning mechanism in the path evaluation module. This mechanism enables the model to adapt to the evolving distribution of explored paths, improving its ability to distinguish plausible from implausible reasoning trajectories. (3) Replacing the context-aware path evaluation module with general LLMs yields degraded performance, confirming the advantage of our fine-tuned path scorer. By capturing fine-grained semantic distinctions among candidate paths, it provides more accurate evaluation signals, enhancing overall search effectiveness.
Method WebQSP CWQ
#Tokens #Calls #Tokens #Calls
DoG 22,538 30.9 37,741 58.1
ToG 16,372 23.2 26,183 41.9
RwT 10,680 15.1 17,885 28.6
DAMR 3,931 7.1 9,266 16.8
Table 2: Statistics of average number of LLM calls and token consumption per question on WebQSP and CWQ datasets.
Method WebQSP CWQ
Hits@1 F1 Hits@1 F1
DAMR w/o PE 91.2 78.2 74.3 72.1
DAMR w/o FT 91.9 80.1 75.1 73.0
DAMR w/ GPT 4.1 92.5 79.8 74.9 72.4
DAMR 94.0 81.7 78.0 75.1
Table 3: The results of ablation studies on the WebQSP and CWQ datasets. Bold results indicate the best performance.
(a) Number of selected relations $k$
(b) Reasoning path length $L$
Figure 2: Sensitivity analysis of hyperparameter on the WebQSP and CWQ datasets.
4.5 Sensitivity Analysis
We conduct a sensitivity analysis to assess the impact of two key hyperparameters in DAMR: the number of selected relations $k$ and the maximum reasoning path length $L$. The parameter $k$ controls how many relations are proposed by the LLM-based planner at each step, while $L$ determines the number of reasoning hops allowed during path construction.
Figure 2 illustrates how $k$ and $L$ affect the performance of DAMR on the WebQSP and CWQ datasets. We vary $k$ and $L$ within the range of $\{2,3,4,5\}$. From the results, we observe that: (1) As shown in Figure 2(a), increasing $k$ initially leads to performance gains, which then stabilize before experiencing a slight decline. While larger $k$ values encourage broader relational exploration, they may also introduce irrelevant candidates and increased computational cost. Conversely, smaller $k$ restrict the diversity of the search. To balance these trade-offs, we select a moderate $k=3$ as the default setting. (2) As shown in Figure 2(b), on the WebQSP dataset, performance improves from $L=2$ to $3$, then fluctuates between $L=3$ and $5$, suggesting limited gains beyond three hops. In contrast, performance on the CWQ dataset steadily increases up to $L=4$ before slightly declining at $L=5$, reflecting its need for deeper reasoning due to more complex questions. Balancing effectiveness and efficiency across both datasets, we set $L=4$ as the default path length in all experiments. More results are provided in Appendix  D.
4.6 Impact of Different LLMs
Method WebQSP CWQ
Hits@1 F1 Hits@1 F1
DAMR (Llama2-13B) 91.0 76.7 73.9 69.5
DAMR (Qwen3-14B) 91.5 77.8 74.4 70.1
DAMR (GPT 4.1-mini) 93.1 80.6 76.1 72.7
DAMR (GPT 4.1) 94.0 81.7 78.0 75.1
Table 4: Performance of DAMR using different LLM-based planners as backbones on different datasets.
To evaluate the impact of different LLM-based planners within the DAMR framework, we compare several backbones including Llama2-13B (roque2025evolution, ), Qwen3-14B (team2024qwen2, ), GPT 4.1 mini, and GPT 4.1, as shown in Table 4. Across both datasets, stronger LLMs consistently yield higher F1 and Hits scores, with GPT 4.1 achieving the best performance across all metrics. These results highlight the pivotal role of advanced LLMs in guiding relation selection and reasoning path expansion, where improved fluency, contextual awareness, and semantic precision translate into more accurate and faithful reasoning trajectories. Notably, the consistent performance gap between smaller and larger backbones illustrates the sensitivity of KGQA systems to the planner’s reasoning capability, reinforcing the necessity of leveraging high-capacity models when available. Overall, these findings emphasize the importance of backbone selection and further validate the design of DAMR, which capitalizes on powerful LLMs to achieve robust, generalizable, and effective multi-hop reasoning.
(a) One example from the WebQSP dataset.
(b) One example from the CWQ dataset.
Figure 3: Case studies of DAMR on the WebQSP and CWQ datasets. We highlight the correct answers in Red and the wrong answers in Blue.
4.7 Case study
Figure 3 presents detailed case studies on the WebQSP and CWQ datasets comparing the reasoning process of DAMR with four representative LLMs: Llama-2-13B, Qwen-3-14B, GPT 4.1-mini, and GPT 4.1. In Figure 3a, the baseline LLMs generate fluent and seemingly plausible answers such as “Proto-Indo-European” or “Proto-Hellenic” when asked about the origin of the Greek language, yet fail to identify the correct answer Attic Group. In contrast, DAMR accurately predicts the correct entity by explicitly traversing the base.rosetta.languoid.parent relation within the knowledge graph. In Figure 3b, none of the baseline models are able to identify the structure that opened in Hollywood in 1922, while the proposed DAMR successfully locates Grauman’s Egyptian Theatre by explicitly following relevant relation paths in the knowledge graph and integrating complementary reasoning trajectories. These examples collectively highlight the strength of DAMR in grounding its reasoning in the knowledge graph and explicitly modeling multi-step reasoning paths, enabling it to provide accurate and faithful answers to complex, ontology-specific questions that general-purpose LLMs often struggle to resolve.
5 Conclusion
In this work, we presented DAMR, a dynamically adaptive MCTS-based framework for knowledge graph question answering that integrates an LLM-guided planner, a context-aware path evaluator, and a dynamic pseudo-path refinement mechanism. By narrowing the search space, reducing redundant LLM calls, and continually adapting path evaluation, DAMR achieves efficient and accurate multi-hop reasoning. Extensive experiments on WebQSP and CWQ demonstrate that DAMR consistently surpasses state-of-the-art baselines in both performance and efficiency, while ablation and case studies highlight its interpretability and robustness. These results establish DAMR as a scalable solution for real-world KGQA and open avenues for extending adaptive symbolic–neural reasoning to broader domains such as scientific discovery and recommendation.
Appendix A Algorithm
Algorithm 1 Dynamic MCTS-based KGQA with Path Model Pretraining and Online Refinement
0: Question $q$, knowledge graph $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, number of selected relations $k$, MCTS iterations NN, length of reasoning path $L$
0: Answer set $\mathcal{A}$
1:/ Stage 1: Path Evaluation Model Pre-training /
2: Construct reasoning path pairs $(q,p^{+},p^{-})$ from $\mathcal{G}$
3: Initialize path evaluation model $S(q,\cdot;\Theta)$
4:for each batch in pretraining data do
5:  Update $S(q,\cdot;\Theta)$ by minimizing the Pair-wise Ranking loss in Eq.(4)
6:end for
7:/ Stage 2: Dynamic MCTS Reasoning /
8:for $i=1$ to NN do
9:  Selection: Traverse the tree from root to a leaf node by selecting child nodes according to the UCT criterion in Eq.(1)
10:  Expansion:
11:  i. At the selected node, enumerate all candidate relations from current entities and use the LLM-based planner to select the top-$k$ most relevant relations
12:  ii. Expand a new child node for each selected relation
13:  Simulation: For each expanded node, perform a rollout by sequentially selecting relations (guided by the path evaluation model) up to $L$ hops or until a correct answer is reached
14:  Backpropagation: Update the value ($w_{i}$) and visit ($n_{i}$) statistics along the traversed path from the leaf node back to the root using the score from the simulation, as per Eq.(6)
15:  Path Evaluation Model Fine-tuning: Generate the explored pseudo-path pairs $(\hat{p}^{+},\hat{p}^{-})$ via Eq.(5) and fine-tune the path evaluation model $S(q,\cdot;\Theta)$ based on Pair-wise Ranking loss in Eq.(4)
16:end for
17:/ Stage 3: Answer Extraction /
18: Collect entities reached by high-scoring reasoning paths as $\mathcal{A}$
19:return $\mathcal{A}$
Appendix B Datasets
B.1 Dataset Description
Datasets #Train #Valid #Test
WebQSP 2,848 250 1,639
CWQ 27,639 3,519 3,531
Table 5: Statistics of KGQA datasets.
We conduct extensive experiments on two widely used multi-hop Knowledge Graph Question Answering (KGQA) benchmarks: WebQSP [47] and CWQ [61]. The statistics of these two benchmarks can be found in Table 5, and their details are shown as follows:
  • The WebQuestionsSP (WebQSP) dataset is a widely adopted benchmark for evaluating single-hop and simple multi-hop KGQA [61]. It consists of 4,837 natural language questions annotated with corresponding SPARQL queries over the Freebase knowledge graph. The dataset is partitioned into 2,848 training, 250 validation, and 1,639 test instances.
  • The ComplexWebQuestions (CWQ) dataset is a challenging benchmark designed for multi-hop KGQA [47]. It comprises 34,689 questions derived from WebQuestionsSP, reformulated to include more complex and compositional queries. Each question typically requires multi-step reasoning over the Freebase knowledge graph, often involving conjunctions, comparatives, or nested logical structures. The dataset is divided into 27,639 training, 3,519 validation, and 3,531 test examples.
B.2 Data Processing
Following prior work [41, 26, 31], we preprocess the datasets by constructing localized subgraphs centered around each question entity to reduce the size of the search space. Specifically, for each question in WebQSP [61] and CWQ [47], we extract a subgraph from the Freebase knowledge graph by including all triples within a predefined number of hops from the topic entity. This approach preserves the essential context required for multi-hop reasoning while significantly improving computational efficiency.
Appendix C Baselines
In this part, we introduce the details of the compared baselines as follows:
  • Semantic Parsing Methods. We compare our DAMR with six semantic parsing methods:
    • KV-Mem: KV-Mem [34] introduce a neural architecture that stores facts as key-value pairs and enables question answering by attending over memory slots, directly retrieving relevant information to infer answers.
    • EmbedKGQA: EmbedKGQA [39] enhances multi-hop question answering over knowledge graphs by leveraging pretrained knowledge base embeddings, enabling the model to reason over entity and relation representations without explicit path enumeration during answer prediction.
    • QGG: QGG [18] generates query graphs to answer multi-hop complex questions over knowledge bases, formulating question answering as query graph prediction and enabling structured reasoning through graph matching and path ranking mechanisms.
    • NSM: NSM [11] enhances multi-hop KBQA by leveraging intermediate supervision signals, decomposing questions into reasoning steps, and training a neural state machine to sequentially predict relations and entities for accurate path-based reasoning.
    • TransferNet: TransferNet [42] proposes a transparent framework for multi-hop QA over relational graphs by transferring question semantics to relation paths through interpretable path ranking and structured reasoning, enabling effective and explainable answer prediction.
    • KGT5: KGT5 [38] formulates knowledge graph completion and question answering as unified sequence-to-sequence tasks, leveraging pre-trained language models to jointly encode input queries and generate answer entities or triples in a flexible and end-to-end manner.
  • Retrieval-Based Methods. We compare our DAMR with four retrieval-based methods:
    • GraftNet: GraftNet [45] proposes an early fusion framework that jointly encodes knowledge base facts and supporting text by constructing a heterogeneous graph, enabling effective reasoning through graph convolutional networks for open-domain question answering.
    • PullNet: PullNet [44] introduces an iterative retrieval mechanism that expands a query-specific subgraph by pulling relevant facts from both knowledge bases and text, enabling joint reasoning over heterogeneous evidence for open-domain question answering.
    • SR+NSM: SR+NSM [64] enhances multi-hop KBQA by first retrieving a question-relevant subgraph and then performing symbolic reasoning over it using Neural Symbolic Machines, improving efficiency and accuracy through constrained and focused logical inference.
    • SR+NSM+E2E: SR+NSM+E2E [64] extends SR+NSM by enabling end-to-end training that jointly optimizes subgraph retrieval and reasoning. This integration enhances model coherence and allows better alignment between retrieved subgraphs and final answer prediction.
  • General Large Language Models (LLMs). We compare our DAMR with six general LLMs:
    • Flan-T5-xl: Flan-T5-xl [5] is an instruction-finetuned variant of the T5 model, trained on a diverse collection of tasks with natural language instructions. By leveraging large-scale instruction tuning, it improves zero-shot and few-shot performance across diverse NLP benchmarks.
    • Alpaca-7B: Alpaca-7B [48] is an instruction-following language model fine-tuned from LLaMA-7B using self-instruct techniques. It demonstrates strong zero-shot and few-shot performance by aligning with human instructions across various NLP tasks.
    • Llama3-8B: Llama3-8B [9] is part of the LLaMA 3 family of models, designed for improved instruction following, reasoning, and code generation. Pretrained on a high-quality corpus and fine-tuned with supervised signals, it achieves strong performance across diverse benchmarks.
    • Qwen2.5-7B: Qwen2.5-7B [49] is a 7B-parameter instruction-tuned language model developed by Alibaba, optimized for tasks such as reasoning, code generation, and dialogue. It supports multi-turn conversation and demonstrates competitive performance on standard benchmarks.
    • ChatGPT: ChatGPT [40] is a conversational AI developed by OpenAI, based on the GPT architecture. It is designed to understand natural language, engage in dialogue, answer questions, and assist with a wide range of tasks across domains.
    • ChatGPT+CoT: ChatGPT with Chain-of-Thought (CoT) [54] prompting enhances the model’s reasoning capabilities by encouraging it to generate intermediate reasoning steps before arriving at a final answer, improving performance on complex, multi-step problems.
  • LLMs with KG. We compare our DAMR with fourteen LLMs with KG methods:
    • UniKGQA: UniKGQA [15] is a unified framework that integrates retrieval and reasoning for multi-hop question answering over knowledge graphs, combining subgraph retrieval, query decomposition, and neural reasoning in an end-to-end manner.
    • DECAF: DECAF [62] is a joint framework for question answering over knowledge bases that simultaneously decodes logical forms and answers. By leveraging dual supervision, it enhances both symbolic reasoning accuracy and direct answer prediction in a unified architecture.
    • KD-CoT: KD-CoT [51] is a framework that enhances the faithfulness of large language models by guiding Chain-of-Thought reasoning with external knowledge, improving accuracy in knowledge-intensive question answering tasks.
    • Nutrea: Nutrea [4] proposes a neural tree search framework for context-guided multi-hop KGQA. It incrementally constructs reasoning trees by integrating question semantics and graph context, enabling efficient exploration of multi-hop paths for accurate answer prediction.
    • ToG: ToG [46] is a framework that enables large language models to perform deep and responsible reasoning over knowledge graphs by combining structured graph information with iterative thinking and verification mechanisms for reliable multi-hop QA.
    • RoG: RoG [28] is a framework that enhances the faithfulness and interpretability of large language model reasoning by grounding multi-hop question answering on knowledge graphs, integrating symbolic path tracking with natural language generation.
    • KAPING: KAPING [1] introduces knowledge-augmented prompting by integrating structured triples into Chain-of-Thought (CoT) reasoning. It guides large language models to generate intermediate reasoning steps, enabling zero-shot multi-hop KGQA without task-specific fine-tuning.
    • ReasoningLM: ReasoningLM [14] enhances pre-trained language models for KGQA by injecting subgraph structures into the input representation. It enables structural reasoning over retrieved subgraphs through a reasoning-aware encoder, improving performance on complex multi-hop queries.
    • FiDeLis: FiDeLis [43] proposes a faithfulness-aware KGQA framework that enhances reasoning consistency in LLMs by aligning generated logical forms with answer predictions. It introduces fidelity constraints to reduce hallucinations and improve answer correctness.
    • GNN-RAG: GNN-RAG [33] integrates graph neural networks with retrieval-augmented generation by encoding knowledge subgraphs into LLMs’ context. It enables structural reasoning over retrieved subgraphs, improving answer accuracy in KGQA through explicit graph-aware representations.
    • DoG: DoG [30] is a flexible and reliable reasoning framework that enables large language models to generate and evaluate multiple reasoning paths over knowledge graphs through a debate-style process, enhancing robustness and answer faithfulness.
    • DuarL: DuarL [23] is a collaborative framework that integrates GNNs and LLMs for KGQA, where GNNs capture structural semantics and LLMs perform adaptive reasoning, enabling accurate and interpretable multi-hop QA.
    • DP: DP [31] is a trustworthy reasoning framework that guides large language models using prior knowledge from knowledge graphs. It iteratively verifies and refines reasoning paths to enhance faithfulness, robustness, and answer accuracy in KGQA.
    • RwT: RwT [41] is a faithful KGQA framework that models multi-hop reasoning as tree-structured exploration over knowledge graphs, enabling large language models to generate interpretable reasoning paths and improve answer consistency and accuracy.
Appendix D More experimental results
To more thoroughly illustrate the impact of hyperparameter variations on model performance, we report detailed numerical results showing how performance fluctuates under different hyperparameter settings. As presented in Table 6 and Table 7, these results provide a comprehensive understanding of the model’s sensitivity and stability across a range of configurations.
Method WebQSP CWQ
Hits@1 F1 Hits@1 F1
$k=2$ 93.0 80.9 76.6 73.8
$k=3$ 94.0 81.7 78.0 75.1
$k=4$ 94.0 81.8 77.8 75.0
$k=5$ 93.9 81.7 78.0 75.2
Table 6: Hyperparameter sensitivity analysis of the number of selected relations $k$ on the WebQSP and CWQ datasets.
Method WebQSP CWQ
Hits@1 F1 Hits@1 F1
$L=2$ 93.6 81.2 77.4 74.5
$L=3$ 94.0 81.7 77.6 74.7
$L=4$ 93.7 81.4 78.0 75.1
$L=5$ 93.8 81.6 77.9 74.9
Table 7: Hyperparameter sensitivity analysis of the reasoning path length $L$ on the WebQSP and CWQ datasets.
Appendix E Prompt Template
We provide the prompt templates used by the LLM-based planner to select the top-$k$ most relevant relations from the candidate set at each step of path expansion in Fig. 3, as part of the LLM Guided Path Expansion module.
Prompt Template for LLM-Guided Path Expansion Role You are an expert assistant for Knowledge Graph Question Answering (KGQA). Your core capability is to deeply understand natural language questions and the semantics of knowledge graph relations to find the most relevant reasoning paths. Task Your task is to act as a “Relation Retriever.” Given a natural language question and a list of candidate relations, you must analyze the semantics of the question and each relation to select up to k relations that are most likely to lead to the correct answer. Rules and Constraints Fidelity to Candidates: Your selection of relations MUST come strictly from the provided Candidate Relations list. Do not invent or modify relations. Quantity Limit: Return no more than k relations. If multiple relations are highly relevant, order them from most to least relevant. If there are fewer than k relevant relations, return only those. Output Format: Your response MUST be a list of strings, containing the names of the relations you have selected. Example Input: Question: "who was the president after jfk died" Candidate Relations: {"government.president", "government.president.successor", "location.location.containedby", "people.person.place_of_birth"} K: 2 Output:
["government.president", "government.president.successor"]
Your Task Question: {question} Candidate Relations: {relations_list} K: {k} Output:
Figure 4: Prompt template used in the LLM-based planner to select top-k relations during reasoning.
Appendix F The Use of Large Language Models (LLMs)
Large language models (LLMs) were only used to improve the clarity, grammar, and fluency of the manuscript. They were not involved in the development of research ideas, experimental design, data analysis, or any other aspect of the scientific content.